OpenStudy (anonymous):

While driving on the freeway at 110 km h, you pass a truck whose total length you estimate at 25 m. If it takes you, in the driver’s seat, 5.5 s to pass from the rear of the truck to its frontHow far does the truck travel while you’re passing it?

OpenStudy (ujjwal):

let the velocity of truck be v your velocity=110 kmph= 30.55 m/s your velocity relative to truck= (30.55-v)m/s you take 5.5 sec to reach from rear to front i.e 25 m so, 25=(30.55-v)\(\times\)5.5=25 Find v from above equation.. Now, Distance traveled by truck while you are passing it i.e. in 5.5 sec is \(v \times 5.5\)